# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ 4.A. Tree of cylinders
Let $G$ be a group acting on a tree $T$. In order to construct an acylindrical action of $G$ on a tree, we will modify the tree $T$ using the technology of tree of cylinders introduced by Guirardel and Levitt [GL11].
@@
Inclusion is an admissible relation for every class of groups ${\mathcal{E}}$. If ${\mathcal{E}}$ is the class of virtually infinite cyclic groups, then commensurability is an admissible equivalence relation, where two groups $A,B\in{\mathcal{E}}$ are commensurable if $A\cap B$ has finite index in both $A$ and $B$.
Let $T$ be an ${\mathcal{E}}$-tree and let $\sim$ be an admissible equivalence relation on ${\mathcal{E}}$. If $e$ is an edge of $T$, we denote by $G_{e}$ its stabiliser in $T$. We define an equivalence relation $\sim_{T}$ on the set of edges of $T$ by setting, for all edges $e,e^{\prime}\in ET$, $e\sim_{T}e^{\prime}$ if and only if $G_{e}\sim G_{e^{\prime}}$. A cylinder $Y$ of $T$ is a $\sim_{T}$-equivalence class, seen as a subforest of $T$. A cylinder is in fact a subtree of $T$ (see [GL11, Lemma 4.2]).
@@
    $V_{0}T_{c}$ is the set of vertices ot $T$ belonging to at least two distinct cylinders;
@@ 4.B. JSJ decompositions of one-ended relatively hyperbolic groups
We now let $G$ be a one-ended hyperbolic group relative to a family ${\mathcal{P}}=\{[P_{1}],\ldots,[P_{n}]\}$ of conjugacy classes of groups and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. Let $G_{\Phi}$ be the suspension $G\rtimes_{\Phi}\mathbb{Z}$. In time, we will also assume the suspensions of the $P_{i}$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, and want to apply Theorem 2.5 to $G_{\Phi}$ in order to prove that $G_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. That is, we will construct a simplicial tree $T$ on which $G_{\Phi}$ acts acylindrically. The construction of the tree $T$ uses the theory of JSJ decomposition of groups, which we now discuss, following the work of Guirardel-Levitt [GL11, GL15, GL17].
A subgroup of $G$ is elementary if it is virtually cyclic or conjugate into some $P_{i}$ with $i\in\{1,\ldots,n\}$. Let ${\mathcal{A}}$ be the family of all elementary subgroups of $G$.
Let $\sim_{\mathcal{A}}$ be the equivalence relation on ${\mathcal{A}}$ given by $A\sim_{\mathcal{A}}B$ if $\langle A,B\rangle$ is elementary. The equivalence relation $\sim_{\mathcal{A}}$ defines an admissible equivalence relation called coelementarity.
Let ${\mathcal{H}}$ be any set of conjugacy classes of subgroups of $G$. Recall that an $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}})$-tree is an ${\mathcal{A}}$-tree $T$ such that, for every $[A]\in{\mathcal{P}}\cup{\mathcal{H}}$, the group $A$ is elliptic in $T$. We denote by $\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}^{(t)})$ the subgroup of $\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}})$ consisting of every $\psi\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}})$ such that, for every $[A]\in{\mathcal{H}}$, there exists $\Psi\in\psi$ with $\Psi(A)=A$ and $\Psi|_{A}=\operatorname{id}_{A}$.
Let $T$ be a tree equipped with an action of $G$ by isometries with a finite number of orbits of edges. If $H$ is a subgroup of $\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}})$ preserving the $G$-equivariant homeomorphism class of a tree $T$, we denote by $H^{0}$ the finite index subgroup of $H$ acting trivially on $G\backslash T$. Note that, for every $v\in VT$, we have a homomorphism $H^{0}\to\operatorname{\mathrm{Out}}(G_{v})$.
Using [GL17, Theorem 9.18], [GL15, Theorem 3.9] and [GL11, Proposition 6.1], we have the following theorem.
@@
[GL11, GL15, GL17] Let $G$ be a one-ended hyperbolic group relative to a family ${\mathcal{P}}$ of non virtually cyclic groups. Let ${\mathcal{H}}$ be any family of conjugacy classes of subgroups of $G$. There exists a tree of cylinders $T_{\mathcal{H}}$ for coelementarity equipped with an isometric action of $G$ such that:
@@
    vertex stabilisers $G_{v}$ corresponding to a vertex $v$ of the original tree satisfy one of the following:
    • $G_{v}$ is nonelementary and Quadratically Hanging (QH) with finite fibre; (see [GL17, Definition 5.13]);
    • the vertex $v$ is nonelementary and rigid: the stabiliser of $v$ is elliptic in every $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}})$-tree (see [GL17, Definition 2.14]).
@@
    if $[H]\in{\mathcal{H}}$ is not elementary, then $H$ stabilises a unique rigid vertex. (This follows from [GL17, Definition 5.13(3)], and the fact that QH vertices with finite fibre have virtually cyclic extended boundary subgroups).
Moreover, if =${\mathcal{H}}=\{[H_{1}],\ldots,[H_{k}]\}$ with every $H_{i}$ finitely generated:
@@
    [GL15, Theorem 3.9] for every rigid vertex $v\in VT_{\mathcal{H}}$, the homomorphism $\operatorname{\mathrm{Out}}^{0}(G,{\mathcal{P}}\cup{\mathcal{H}}^{(t)})\to\operatorname{\mathrm{Out}}(G_{v})$ is finite;
@@
When the family ${\mathcal{H}}$ is trivial, we will refer to $T_{\mathcal{H}}$ as $T^{\mathrm{can}}$. (Our superscript convention here is certainly not standard: we use it because we will shortly need to discuss minimal invariant trees for subgroups coming from multiple underlying actions. This choice lets us write $T^{\mathrm{can}}_{H}$, for instance, keeping both the tree and the subgroup conveniently in the notation.)
We now prove a general lemma in order to deduce acylindrical actions of $G_{\Phi}$ with $\Phi\in\operatorname{\mathrm{Aut}}(G)$ on trees out of acylindrical actions of $G$. If $G_{\Phi}$ acts on a tree $T$, we denote by $F_{\Phi}$ the $G$-equivariant isometry of $T$ induced by $\Phi$.
@@
Let $K\geq 1$, and let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G)$. Suppose that $G_{\Phi}$ acts on a tree $T$ with finitely many orbits of edges and that the action of $G$ on $T$ is $K$-acylindrical. Suppose that for every geodesic path $\gamma$ of length $3$ and every automorphism $\Psi\in\phi$ such that $F_{\Psi}$ preserves $\gamma$, there exist a vertex $v$ of $\gamma$ and $g\in G_{v}$ of infinite order fixed by a power of $\Psi$.
@@
    Let $n\in\mathbb{N}^{*}$ and let $\Psi\in\phi^{n}$. Suppose that $F_{\Psi}$ fixes pointwise a geodesic edge path of length at least equal to $72K+$ 7. There exists $N\in\mathbb{N}^{*}$ such that $\Psi^{N}$ fixes elementwise a nonabelian free group $L\subseteq G$ consisting of loxodromic elements of $T$. Moreover, for every $g\in L$, the isometry $F_{\Psi^{N}}$ fixes elementwise the axis of $g$.
  2. (2)
    Let $n\in\mathbb{N}^{*}$ and let $\Psi\in\phi^{n}$. Suppose that there exists $g\in\operatorname{\mathrm{Fix}}(\Psi)$ which acts loxodromically on $T$. There exist $m\in\mathbb{Z}$, $n\in\mathbb{Z}^{*}$ such that $g^{m}F_{\Psi^{n}}$ fixes pointwise the axis of $g$.
@@
$(1)$ Suppose that $F_{\Psi}$ fixes pointwise a geodesic edge path $\gamma$ of length $72K+$ 7. Thus, the path $\gamma$ is not reduced to an edge and the isometry $F_{\Psi}$ fixes the initial and the terminal paths $\gamma_{1},\gamma_{2}$ of $\gamma$ length $3$. Since the action of $G$ on $T$ is $K$-acylindrical, for all vertices $v_{1}$ of $\gamma_{1}$ and $v_{2}$ of $\gamma_{2}$, the intersection $G_{v_{1}}\cap G_{v_{2}}$ is finite.
@@
$(2)$ Let $g\in\operatorname{\mathrm{Fix}}(\Psi)$ be a loxodromic element. As in the proof of Assertion $(1)$, the axis of $g$ is preserved by $F_{\Psi}$ and is contained in the characteristic set of $F_{\Psi}$. Thus, we have a homomorphism $\Lambda\colon\langle g,F_{\Psi}\rangle\to\mathbb{R}$ given by the translation length on the axis of $g$. Since $T$ is a simplicial tree, the image of $\Lambda$ is a discrete subset of $\mathbb{R}$. Thus, the image of $\Lambda$ is cyclic. Therefore, the kernel of $\Lambda$ is nontrivial: there exist $m\in\mathbb{Z}$ and $n\in\mathbb{Z}^{*}$ such that $g^{m}F_{\Psi^{n}}$ fixes pointwise the axis of $g$. This proves Assertion $(2)$.
$(3)$ Suppose that there exist $n\in\mathbb{N}^{*}$, $\Psi\in\phi^{n}$ and $g\in\operatorname{\mathrm{Fix}}(\Psi)$ such that the action of $g$ on $T$ is loxodromic. By Assertion $(2)$, there exist $m\in\mathbb{Z}$ and $n\in\mathbb{Z}^{*}$ such that $g^{m}T_{\Psi^{n}}$ fixes pointwise the axis of $g$. Hence the action of $G_{\Phi}$ is not acylindrical.
Conversely, suppose that the action of $G_{\Phi}=\langle G,t\rangle$ on $T$ is not acylindrical. Since the action of $G$ on $T$ is $K$-acylindrical, there exist $g\in G$ and $k\in\mathbb{N}^{*}$ such that the element $gt^{k}$ fixes an edge path of length $72K+$ 7. Let $\Psi\in\phi^{k}$ be the automorphism corresponding to $gt^{k}$. By Assertion $(1)$, some power of $\Psi$ fixes a loxodromic element of $G$. This proves Assertion $(3)$ and concludes the proof. ∎
@@ 5. The periodic JSJ decomposition
We now specialise to the tree we will use to prove that suspensions of one-ended relatively hyperbolic groups (under reasonable assumptions on the parabolic subgroups) satisfy the Farrell–Jones conjecture. Let $G$ be a one-ended relatively hyperbolic group and let $\Phi\in\operatorname{\mathrm{Aut}}(G)$. We explain in the following section the construction of $G$-trees which are naturally associated with $\Phi$.

### 02 Mar 2026 v2

@@ 4.A. Tree of cylinders
Let $G$ be a group acting on a tree $T$. In order to construct an acylindrical action of $G$ on a tree, we will modify the tree $T$ using the technology of tree of cylinders introduced by Guirardel and Levitt [26].
@@
The equivalence relation generated by inclusion is an admissible relation for every class of groups $\mathcal{E}$} (see [26, Lemma 3.8]). If $\mathcal{E}$} is the class of virtually infinite cyclic groups, then commensurability is an admissible equivalence relation, where two groups $A,B\in{\mathcal{E}}$ are commensurable if $A\cap B$ has finite index in both $A$ and $B$.
Let $T$ be an $\mathcal{E}$}-tree and let $\sim$ be an admissible equivalence relation on $\mathcal{E}$}. If $e$ is an edge of $T$, we denote by $G_{e}$ its stabiliser in $T$. We define an equivalence relation $\sim_{T}$ on the set of edges of $T$ by setting, for all edges $e,e^{\prime}\in ET$, $e\sim_{T}e^{\prime}$ if and only if $G_{e}\sim G_{e^{\prime}}$. A cylinder $Y$ of $T$ is a $\sim_{T}$-equivalence class, seen as a subforest of $T$. A cylinder is in fact a subtree of $T$ (see [26, Lemma 4.2]).
@@
    $V_{0}T_{c}$ is the set of vertices of $T$ belonging to at least two distinct cylinders;
@@ 4.B. JSJ decompositions of one-ended relatively hyperbolic groups
We now let $G$ be a hyperbolic group relative to a family ${\mathcal{P}}=\{[P_{1}],\ldots,[P_{n}]\}$ of conjugacy classes of groups. Let $\mathcal{H}$} be a set of conjugacy classes of subgroups of $G$ such that $G$ is one-ended relative to $\mathcal{P}$$\cup{\mathcal{H}}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}}\cup{\mathcal{H}})$. Let $G_{\Phi}$ be the suspension $G\rtimes_{\Phi}\mathbb{Z}$. In time, we will also assume the suspensions of the $P_{i}$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, and want to apply Theorem 2.5 to $G_{\Phi}$ in order to prove that $G_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. That is, we will construct a simplicial tree $T$ on which $G_{\Phi}$ acts acylindrically. The construction of the tree $T$ uses the theory of JSJ decomposition of groups, which we now discuss, following the work of Guirardel-Levitt [26, 27, 30].
A subgroup of $G$ is elementary if it is finite, or virtually cyclic or conjugate into some $P_{i}$ with $i\in\{1,\ldots,n\}$. Let $\mathcal{A}$} be the family of all infinite elementary subgroups of $G$.
Let $\sim_{\mathcal{A}}$ be the equivalence relation on $\mathcal{A}$} given by $A\sim_{\mathcal{A}}B$ if $\langle A,B\rangle$ is elementary. The equivalence relation $\sim_{\mathcal{A}}$ defines an admissible equivalence relation called coelementarity (see [26, Lemma 3.4]).
Let $\mathcal{H}$} be any set of conjugacy classes of subgroups of $G$. Recall that an $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}})$-tree is an $\mathcal{A}$}-tree $T$ such that, for every $A]\in{\mathcal{P}}\cup{\mathcal{H}}$, the group $A$ is elliptic in $T$. We denote by $\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}^{(t)})$ the subgroup of $\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}})$ consisting of every $\psi\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}})$ such that, for every $A]\in{\mathcal{H}}$, there exists $\Psi\in\psi$ with $\Psi(A)=A$ and $\Psi|_{A}=\operatorname{id}_{A}$.
[…]
Using [30, Theorem 9.18], [27, Theorem 3.9] and [26, Proposition 6.1], we have the following theorem.
@@
[26, 27, 30] Let $G$ be a hyperbolic group relative to a family $\mathcal{P}$} of non virtually cyclic groups. Let $\mathcal{H}$} be any family of conjugacy classes of subgroups of $G$. Suppose that $G$ is one-ended relative to $\mathcal{P}$$\cup{\mathcal{H}}$. There exists a tree of cylinders $T_{\mathcal{H}}$ for coelementarity equipped with an isometric action of $G$ such that:
@@
    the vertex stabiliser $G_{v}$ of a noncylinder vertex $v$ satisfies one of the following:
    • $G_{v}$ is nonelementary and Quadratically Hanging (QH) with finite fibre; (see [30, Definition 5.13]);
    • the vertex $v$ is nonelementary and rigid: the stabiliser of $v$ is elliptic in every $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}})$-tree (see [30, Definition 2.14]).
@@
    if $H]\in{\mathcal{H}}$ is not elementary, then $H$ stabilises a unique rigid vertex. (Uniqueness follows from $(4)$. The fact that the vertex is rigid follows from [30, Definition 5.13(3)], and the fact that QH vertices with finite fibre have virtually cyclic extended boundary subgroups).
Moreover, if ${\mathcal{H}}=\{[H_{1}],\ldots,[H_{k}]\}$ with every $H_{i}$ finitely generated:
@@
    [27, Theorem 3.9] for every rigid vertex $v\in VT_{\mathcal{H}}$, the homomorphism $\operatorname{\mathrm{Out}}^{0}(G,{\mathcal{P}}\cup{\mathcal{H}}^{(t)})\to\operatorname{\mathrm{Out}}(G_{v})$ is finite;
@@
When the family $\mathcal{H}$} is explicit from context, we will refer to $T_{\mathcal{H}}$ as $T^{\mathrm{can}}$. (Our superscript convention here is certainly not standard: we use it because we will shortly need to discuss minimal invariant trees for subgroups coming from multiple underlying actions. This choice lets us write $T^{\mathrm{can}}_{H}$, for instance, keeping both the tree and the subgroup conveniently in the notation.)
We now prove a sequence of lemmas in order to deduce acylindrical actions of $G_{\Phi}$ with $\Phi\in\operatorname{\mathrm{Aut}}(G)$ on trees out of acylindrical actions of $G$. If $G_{\Phi}$ acts on a tree $T$, we denote by $F_{\Phi}$ the $G$-equivariant isometry of $T$ induced by $\Phi$.
Whilst we are certain the next lemma is known to experts we did not manage to locate it in the literature and so provide a proof.
@@
If a group $G$ acts acylindrically on a tree $T$, then every subgroup $H\leqslant G$ also acts acylindrically on $T$. Moreover, if $H$ has no loxodromic elements, then $H$ has a global fixed point on $T$.
Proof.
The first assertion is easy, since if for every $(K_{1},K_{2})\in\mathbb{N}^{2}$ there is a path of length at least $K_{1}$ stabilised by a subgroup of $H$ of size at least $K_{2}$ then the same is true for $G$. For the second assertion, observe that a subgroup all of whose elements are elliptic but which does not have a global fixed point will be a proper ascending union $H_{1}<H_{2}<\cdots<H_{n}<\cdots$, where each $H_{i}$ stabilises an infinite ray. For each $K_{2}$ there will be a subgroup $H_{i}$ with greater cardinality, and so such an action cannot be acylindrical. ∎
Lemma 4.4.
Suppose $T$ is a $G$-tree, preserved by $\phi\in\operatorname{\mathrm{Out}}(G)$, and for $\Phi\in\phi$ let $F_{\Phi}$ be the induced $G$-equivariant isometry of $T$. If $g$ is loxodromic and contained in $\operatorname{\mathrm{Fix}}(\Phi)$, then there are $m\in\mathbb{Z}$ and $n\in\mathbb{Z}^{*}$ so that $g^{m}F_{\Phi}^{n}$ fixes pointwise the axis of $g$.
Note that it could be true that $\Phi^{n}$ or $F_{\Phi^{n}}$ is trivial, even if $n$ is not.
Proof.
Since $g$ is contained in $\operatorname{\mathrm{Fix}}(\Phi)$, the isometry $F_{\Phi}$ commutes with $g$ and must preserve the axis of $g$ and its orientation. This implies that the axis is contained in the characteristic set of $F_{\Phi}$. Thus, we have a homomorphism $\Lambda\colon\mathbb{Z}^{2}\to\langle g,F_{\Phi}\rangle\to\mathbb{R}$ given by the translation length on the axis of $g$. Since $T$ is a simplicial tree, the image of $\Lambda$ is a discrete subset of $\mathbb{R}$. Thus, the image of $\Lambda$ is cyclic and the kernel of $\Lambda$ is nontrivial. Since it cannot be contained in $\langle g\rangle$, there are $m\in\mathbb{Z}$ and $n\in\mathbb{Z}^{*}$ such that $g^{m}F_{\Phi^{n}}$ fixes pointwise the axis of $g$. ∎
Lemma 4.5.
Let $K\geq 1$, and let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G)$. Suppose that $G_{\Phi}$ acts on a tree $T$ with finitely many orbits of edges and that the action of $G$ on $T$ is $K$-acylindrical. Suppose that there exists $J\in\mathbb{N}^{*}$ so that for every geodesic path $\gamma$ of length $J$ and every automorphism $\Psi\in\phi$ such that $F_{\Psi}$ preserves $\gamma$, there exist a vertex $v$ of $\gamma$ and $g\in G_{v}$ of infinite order fixed by a power of $\Psi$.
@@
    Let $n\in\mathbb{N}^{*}$ and let $\Psi\in\phi^{n}$. Suppose that $F_{\Psi}$ fixes pointwise a geodesic edge path of length at least equal to $K+2J+1$. There exists $N\in\mathbb{N}^{*}$ such that $\Psi^{N}$ fixes elementwise a nonabelian free group $L\subseteq G$ consisting of loxodromic elements of $T$. Moreover, for every $g\in L$, the isometry $F_{\Psi^{N}}$ fixes elementwise the axis of $g$.
@@
$(1)$ Suppose that $F_{\Psi}$ fixes pointwise a geodesic edge path $\gamma$ of length $K+2J+1$. Thus, the path $\gamma$ is not reduced to an edge and the isometry $F_{\Psi}$ fixes the initial and the terminal paths $\gamma_{1},\gamma_{2}$ of $\gamma$ length $J$. Since the action of $G$ on $T$ is $K$-acylindrical, for all vertices $v_{1}$ of $\gamma_{1}$ and $v_{2}$ of $\gamma_{2}$, the intersection $G_{v_{1}}\cap G_{v_{2}}$ is finite.
@@
$(2)$ By Lemma 4.3, $\operatorname{\mathrm{Fix}}(\Psi)$ is either elliptic or contains an element acting loxodromically. Suppose that there exist $n\in\mathbb{N}^{*}$, $\Psi\in\phi^{n}$ and $g\in\operatorname{\mathrm{Fix}}(\Psi)$ such that the action of $g$ on $T$ is loxodromic. By Lemma 4.4, there exist $m\in\mathbb{Z}$ and $n\in\mathbb{Z}^{*}$ such that $g^{m}F_{\Psi^{n}}$ fixes pointwise the axis of $g$. Hence the action of $G_{\Phi}$ is not acylindrical.
Conversely, suppose that the action of $G_{\Phi}=\langle G,t\rangle$ on $T$ is not acylindrical. Since the action of $G$ on $T$ is $K$-acylindrical, there exist $g\in G$ and $k\in\mathbb{N}^{*}$ such that the element $gt^{k}$ fixes an edge path of length $K+2J+1$. Let $\Psi\in\phi^{k}$ be the automorphism corresponding to $gt^{k}$. By Assertion $(1)$, some power of $\Psi$ fixes a loxodromic element of $G$. This proves Assertion $(2)$ and concludes the proof. ∎
@@ 5. The periodic JSJ decomposition
We now specialise to the tree we will use to prove that suspensions of one-ended relatively hyperbolic groups (under reasonable assumptions on the parabolic subgroups) satisfy the Farrell–Jones conjecture. Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that $G$ is one-ended relative to $\mathcal{P}\cup{\mathcal{H}}_{0}$. Let $\Phi\in\operatorname{\mathrm{Aut}}(G,\mathcal{P}\cup{\mathcal{H}}_{0}^{(t)})$. We explain in the following section the construction of $G$-trees which are naturally associated with $\Phi$. Note that for the results needed for the Farrell–Jones conjecture, we take ${\mathcal{H}}_{0}$ to be empty; but we will need it to be non-empty for an argument in Section˜5.C.